# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ 4.1 Constrained Multi-Objective Gradient Aggregator
$\mathrm{min}_{g\in\Theta}\;||g||_{H}\;\mathrm{s.t.}\;e_{i}\leq J_{R_{i}}(\theta_{\mathrm{old}}+g)-J_{R_{i}}(\theta_{\mathrm{old}}),$ (6)
@@
where $e_{i}:=\epsilon||g_{i}||_{H^{-1}}$. It is shown that the solutions to (5) and (6) are identical in Appendix A.1, and this finding confirms that objectives can be converted into constraints in this way. Before extending this process to the original problem (2), we scale $e_{i}$ in (6), which represents the degree of increase in $J_{R_{i}}(\theta)$, by the preference value $\omega_{i}$ to reflect a given preference $\omega$. Consequently, the transformed problem is written as follows:
[…]
$J_{C_{k}}(\theta_{\mathrm{old}}+g)\leq d_{k}\;\forall k.$
Now, we calculate the policy gradient by solving (7) through approximation. By linearly approximating the objective and constraint functions, (7) becomes a quadratic programming (QP) problem, which can be solved using a QP solver (Goldfarb & Idnani, 1983). However, due to the approximation, there is a risk of violating the constraints during a policy update. In order to handle this issue, we marginally reduce the threshold to $d_{k}^{\prime}:=d_{k}-K\epsilon$, where $K\in\mathbb{R}_{\geq 0}$ is a hyperparameter. Additionally, if the constraint values are significantly lower than the thresholds, there is no need to consider those constraints. Therefore, we only consider active constraints, whose indices are defined as $I^{\epsilon}_{\mathrm{AC}}:=\{k|d_{k}^{\prime}-\epsilon||b_{k}||_{H^{-1}}\leq J_{C_{k}}(\theta_{\mathrm{old}})\}$. Then, the aggregated gradient can be calculated by solving the following problem using a QP solver:
@@
$\bar{g}_{\omega}^{\mathrm{ag}}:=\mathrm{argmin}_{g}||g||_{H}\;\mathrm{s.t.}\;\omega_{i}e_{i}\leq g_{i}^{T}g\;\forall i,$ (8)
$b_{k}^{T}g+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}^{\prime}\;\forall k\in I_{\mathrm{AC}}^{\epsilon}.$
For more details on solving (8), see Appendix B.3.2. Since $e_{i}$ is proportional to the scale of $g_{i}$, only the directional component of the gradient $g_{i}$ remains in the transformed constraints. Due to this, the proposed method is not affected by the scale of objectives. Furthermore, a policy updated using the aggregated gradient with a proper step size can enhance the objectives and satisfy the constraints.
Lemma 4.1.
Assume that the gradients of the objective and constraint functions are $L$-Lipschitz continuous. If $J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}$ for $k\in I_{\mathrm{AC}}^{\epsilon}$, updating the policy parameter with the following step size $\beta$ ensures $J_{R_{i}}(\theta_{\mathrm{old}})\leq J_{R_{i}}(\theta_{\mathrm{new}})$ for $i\in I_{\omega}$ and $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$ for $k\in I_{\mathrm{AC}}^{\epsilon}$:
[…]
$\beta\leq 2||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}/(\alpha L||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2}),$
$\alpha:=||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}/\mathrm{min}(K\epsilon,\mathrm{min}_{i\in I_{\omega}}(e_{i}))$.
The proof is provided in Appendix A.2. Lemma 4.1 shows that if the step size satisfies (9), updating a policy with $\bar{g}_{\omega}^{\mathrm{ag}}$ not only improves the objectives but also consistently keeps the active constraints below the thresholds. We now propose a proper $\beta$ for stable policy training while meeting the above conditions. From (9), it is observed that $\beta$ should be reduced when the magnitude of $\bar{g}_{\omega}^{\mathrm{ag}}$ and $\alpha$ increase. Through this observation, we suggest clipping the policy gradient such that $\bar{g}_{\omega}^{\mathrm{ag}}$ does not exceed the local region. Similarly, we scale down the policy gradient if $\alpha$ exceeds a specific value, denoted by $G\in\mathbb{R}_{>0}$. Therefore, the final proposed step size is expressed as follows:
$g_{\omega}^{\mathrm{ag}}:=\beta_{1}\beta_{2}\bar{g}_{\omega}^{\mathrm{ag}},\;\text{where}\;$ $\beta_{1}:=\mathrm{min}(1,\epsilon/||\bar{g}_{\omega}^{\mathrm{ag}}||_{H}),$ (10)
$\beta_{2}:=\mathrm{min}(1,G/\alpha),$
where $G$ is a hyperparameter. Then, we can show that the updated policy using the above gradient can improve the objectives and maintain constraint satisfaction.
@@
Assume that the gradients of the objectives and constraints are $L$-Lipschitz continuous. If $J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}$ $\forall k$, there exists $\epsilon>0$ such that the updated policy parameter using CoMOGA, $\theta_{\mathrm{new}}=\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}$, improves $J_{R_{i}}(\theta)$ for $i\in I_{\omega}$ and satisfies $J_{C_{k}}(\theta_{\mathrm{new}})\leq d_{k}$ $\forall k$.
The proof is provided in Appendix A.2 and shows that a larger $G$ or a smaller $K$ necessitates a smaller $\epsilon$. To better understand the role of each hyperparameter, $\epsilon$, $K$, and $G$, please refer to the results of ablation studies in Appendix C. Finally, the process of CoMOGA is illustrated in Figure 2.
4.2 Universal Policy Update
We have introduced a gradient aggregation method named CoMOGA for updating a policy for a given single preference. This section presents a method for updating a universal policy that can cover the entire preference space. To achieve this, we randomly select $P$ preferences, $\{\omega^{p}\}_{p=1}^{P}$, from the preference space and calculate the corresponding aggregated gradients for each preference. Using these $P$ gradients, we can formulate $P$ intermediate policies, defined as $\bar{\pi}_{p}:=\pi_{\theta_{p}}$, where $\theta_{p}:=\theta_{\mathrm{old}}+g^{\mathrm{ag}}_{\omega^{p}}$. Now, we need to find a parameter of the universal policy such that $\pi_{\theta_{\mathrm{new}}}(\cdot|\cdot,\omega^{p})=\bar{\pi}_{p}(\cdot|\cdot,\omega^{p})$ for $\forall p=1,2,...,P$. To this end, the final step involves updating the parameter to minimize the KL divergence across all the intermediate policies as follows:
[…]
where $\mathcal{D}$ is a replay buffer. By minimizing the above loss, we can combine policies for each preference into a single universal policy.
4.3 Convergence Analysis
This section analyzes whether the proposed method converges and, if so, where it converges. Our analysis is grounded on the following three assumptions:
Assumption 4.3.
1) The gradients of the objectives and constraints are $L$-Lipschitz continuous, 2) the constraints are convex, 3) and $\{\theta|J_{C_{k}}(\theta)<d_{k}^{\prime},\;\forall k\}\neq\emptyset$.
The first assumption is common in multi-objective approaches for multi-task learning (Liu et al., 2021; Navon et al., 2022). The last two assumptions are used in multi-constrained RL works (Kim et al., 2023) and are intended to ensure the existence of a solution that satisfies the constraints, which can be viewed as the Slater condition (Boyd & Vandenberghe, 2004). Before proceeding further, we define a concept termed constrained-Pareto (CP) stationarity.
Definition 4.4 (Constrained-Pareto Stationarity).
Given a feasible policy parameter $\theta\in\{\theta|J_{C_{k}}(\theta)\leq d_{k}\;\forall k\}$, the policy $\pi_{\theta}$ is CP stationary if there exist $\lambda\in\mathbb{R}^{N}_{\geq 0}$ and $\nu\in\mathbb{R}^{M}_{\geq 0}$ that satisfy the following:
$\small\sum\nolimits_{i=1}^{N}\lambda_{i}\nabla_{\theta}J_{R_{i}}(\theta)=\sum\nolimits_{k=1}^{M}\nu_{k}\nabla_{\theta}J_{C_{k}}(\theta),$ (12)
where $\sum_{i}\lambda_{i}=1$ and $\nu_{k}=0$ for $k\in\{k|J_{C_{k}}(\theta)<d_{k}\}$.
We add constraint-related terms to the definition of the Pareto stationarity (Désidéri, 2012). We can then show that CP stationarity is a necessary condition for CP optimality.
@@ 4.3 Convergence Analysis
If a policy $\pi_{\theta}$ is CP optimal, the policy is CP stationary. If the objectives are concave, the opposite also holds true.
The proof can be found in Appendix A.2. Through Theorem 4.5, a CP stationary policy can be viewed as a local CP optimal policy. Additionally, we further define the following relaxed stationarity condition.
Definition 4.6 ($\epsilon$-CP Stationarity).
For a policy parameter $\theta\in\{\theta|J_{C_{k}}(\theta)\leq d_{k}\;\forall k\}$, the policy is $\epsilon$-CP stationary if there exist $\lambda\in\mathbb{R}^{N}_{\geq 0}$ and $\nu\in\mathbb{R}^{M}_{\geq 0}$ satisfying (12), $\sum_{i}\lambda_{i}=1$, and $\nu_{k}=0$ for $k\notin I_{\mathrm{AC}}^{\epsilon}$, which is defined in (8).
The definition is the same as CP stationarity except the condition for $\nu_{k}=0$. When $\epsilon=0$, $\epsilon$-CP stationarity is identical to CP stationarity, so the CP stationary set is a subset of the $\epsilon$-CP stationary set. We can now show that training a policy using CoMOGA converges to a $\epsilon$-CP stationary policy.
Theorem 4.7.
By updating a policy using CoMOGA, the policy converges to a $\epsilon$-CP stationary policy. Furthermore, if there is no solution to (8), the current policy is $\epsilon$-CP stationary.
The proof of Theorem 4.7 is presented in Appendix A.2. For a simple explanation of the proof, the proposed method monotonically improves the objectives for $i\in I_{\omega}$. Then, the updated policy will reach some point where either the objectives do not increase or a solution for (8) does not exist. At this point, we have shown the existence of $\lambda$ and $\nu$ satisfying the conditions of Definition 4.6.
@@ 5 Practical Implementation
This section covers practical considerations related to algorithm implementation, such as calculating the gradients of the objectives and handling feasibility issues.
Gradient Calculation. To perform CoMOGA and the universal policy update, we need to calculate the gradient of the objective and constraint functions. To do that, we estimate the objectives and constraints using reward and cost critics $V_{R,\psi_{R}}^{\pi}$, $V_{C,\psi_{C}}^{\pi}$, where $\psi_{R}$ and $\psi_{C}$ are the critic network parameters, for a given preference $\omega$ as follows:
$J_{R_{i}}(\theta)$ $\approx\mathbb{E}_{s\sim\mathcal{D}}\left[V_{R_{i},\psi_{R}}^{\pi}(s,a,\omega)|a\sim\pi_{\theta}(\cdot|s,\omega)\right],$ (13)
$J_{C_{k}}(\theta)$ $\approx\mathbb{E}_{s\sim\mathcal{D}}\left[V_{C_{k},\psi_{C}}^{\pi}(s,a,\omega)|a\sim\pi_{\theta}(\cdot|s,\omega)\right].$
Then, we can compute the gradients using the critic networks as done in the SAC algorithm (Haarnoja et al., 2018).
Feasibility Handling. CoMOGA ensures that if the current policy satisfies the constraints, the updated policy also satisfies them. However, if the current policy does not satisfy the constraints, the policy should be recovered to the feasible region, $\{\theta|J_{C_{k}}(\theta)\leq d_{k}\;\forall k\}$. To handle this issue, we utilize the gradient integration method (GIM) proposed by Kim et al. (2023), which calculates a gradient toward the feasible region under multiple constraint settings. Finally, the proposed method is summarized in Algorithm 1.
Algorithm 1 Policy Update Rule Using CoMOGA
@@
       Collect a trajectory $\tau=\{(s_{t},a_{t},r_{t},c_{t},s_{t+1})\}_{t=1}^{T}$ by using $\pi_{\theta}(\cdot|\cdot,\omega)$ and store $\tau$ in $\mathcal{D}$.
@@
    Update reward and cost critic networks with $\mathcal{D}$.
@@
       Sample $\omega^{p}\sim\Omega$, and estimate the objectives and constraints using Equation (13).
       if Constraints are satisfied then
          Calculate the aggregated gradient $g^{\mathrm{ag}}_{\omega^{p}}$ using Equation (8), and set $\theta_{p}=\theta+g^{\mathrm{ag}}_{\omega^{p}}$.
       else
          Calculate the recovery gradient $g_{r}$ using GIM as described in Section 5, and set $\theta_{p}=\theta+g_{r}$.
       end if
       Store an intermediate policy $\bar{\pi}_{p}=\pi_{\theta_{p}}$.

### v2

@@ 4.1 Transformation to Constrained Optimization
$\mathrm{min}_{\Delta\theta}\;\Delta\theta^{T}H\Delta\theta\quad\mathbf{s.t.}\;e_{i}\leq J_{R_{i}}(\theta_{\mathrm{old}}+\Delta\theta)-J_{R_{i}}(\theta_{\mathrm{old}}),$ (3)
@@
where $e_{i}:=\epsilon||g_{i}||_{H^{-1}}$. It is shown that the solutions to (2) and (3) are identical in Appendix B.1, and this finding confirms that objectives can be converted into constraints in this way. Consequently, the CMORL problem can be transformed as follows:
$\mathrm{min}_{\Delta\theta}\Delta\theta^{T}H\Delta\theta\quad\mathbf{s.t.}\;\omega_{i}e_{i}\leq J_{R_{i}}(\theta_{\mathrm{old}}+\Delta\theta)-J_{R_{i}}(\theta_{\mathrm{old}})\;\forall i,\;J_{C_{k}}(\theta_{\mathrm{old}}+\Delta\theta)\leq d_{k}\;\forall k,$ (4)
where $e_{i}$ in (3) is scaled to $\omega_{i}e_{i}$ to reflect the given preference $\omega$. We will show that the policy updated using (4) converges to a CP optimal policy of (1) in Section 4.4 under mild assumptions.
4.2 Gradient Aggregation
Now, we calculate a policy gradient by solving (4) through a linear approximation and update a policy within the local region to reduce approximation errors. First, the problem (4) can be linearly approximated as the following quadratic programming (QP) problem:
@@ 4.2 Gradient Aggregation
$\bar{g}_{\omega}^{\mathrm{ag}}:=\mathrm{argmin}_{\Delta\theta}\Delta\theta^{T}H\Delta\theta\quad\mathbf{s.t.}\;\omega_{i}e_{i}\leq g_{i}^{T}\Delta\theta\;\forall i,\;b_{k}^{T}\Delta\theta+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}\;\forall k,$ (5)
where the gradients of the objective and constraint functions, $g_{i}$ and $b_{k}$, are aggregated into $\bar{g}_{\omega}^{\mathrm{ag}}$. The gradient is then clipped to ensure that the policy is updated within the local region, as follows:
$g_{\omega}^{\mathrm{ag}}:=\min(1,\epsilon/||\bar{g}_{\omega}^{\mathrm{ag}}||_{H})\bar{g}_{\omega}^{\mathrm{ag}},$
where the policy will be updated by $\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}$. However, due to the linear approximation, the updated policy can violate the safety constraints. To address this issue, we take a recovery step by solving the following QP problem, which only consists of the safety constraints, when the current policy violates any safety constraints:
$\bar{g}_{\omega}^{\mathrm{ag}}:=\mathrm{argmin}_{\Delta\theta}\Delta\theta^{T}H\Delta\theta\quad\mathbf{s.t.}\;b_{k}^{T}\Delta\theta+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}\;\forall k.$ (6)
As a result, the process for obtaining the policy gradient can be summarized as follows, which is named constrained multi-objective gradient aggregator (CoMOGA):
$\bar{g}_{\omega}^{\mathrm{ag}}$ =$=\begin{cases}\mathrm{argmin}_{\Delta\theta}\Delta\theta^{T}H\Delta\theta\;\mathbf{s.t.}\;\omega_{i}e_{i}\leq g_{i}^{T}\Delta\theta,\;b_{k}^{T}\Delta\theta+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}&\text{if}\;J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}\;\forall k,\\ \mathrm{argmin}_{\Delta\theta}\Delta\theta^{T}H\Delta\theta\;\mathbf{s.t.}\;b_{k}^{T}\Delta\theta+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}&\text{otherwise,}\\ \end{cases}$
$g_{\omega}^{\mathrm{ag}}$ $=\min(1,\epsilon/||\bar{g}_{\omega}^{\mathrm{ag}}||_{H})\bar{g}_{\omega}^{\mathrm{ag}},$
and it is illustrated in Figure 3. Note that the aggregated gradient satisfies $0\leq g_{i}^{T}g_{\omega}^{\mathrm{ag}}$ $\forall i$ due to the transformed constraints in (5), which ensures that the aggregated gradient does not conflict with any objective functions.
Figure 3: Process of CoMOGA. We visualize the process in the parameter space, and the gray areas represent constraints. (Linear approximation) CoMOGA linearly approximates the original CMORL problem in (1). The gradients of the objective and constraint functions are visualized as black and blue arrows, respectively. (Transformation) The objectives are converted to constraints as described in (5). The intersection of all constraints is shown as the red area. (Scaling) The solution of the transformed problem, $\bar{g}^{\mathrm{ag}}$, is scaled to ensure that the updated policy remains within the local region.
4.3 Update Rule for Universal Policy
We have introduced a gradient aggregation method named CoMOGA for updating a policy for a given single preference. This section presents a method for updating a universal policy that can cover the entire preference space. Given that a policy is updated to $\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}$ when a preference $\omega$ is specified, the ideal universal policy should satisfy the following conditions:
$\bar{\pi}(a|s,\omega)=\pi_{\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}}(a|s,\omega)\quad\forall a\in A,\;\forall s\in S,\;\text{and}\;\forall\omega\in\Omega,$ (7)
where we denote $\pi_{\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}}$ as an intermediate policy $\bar{\pi}_{\omega}$. This can be interpreted as the KL divergence between the universal policy and the intermediate policy being zero. Based on this property, we introduce a practical method to achieve the universal policy by minimizing the following loss:
$\mathrm{min}_{\theta}\;\mathbb{E}_{(s,\omega)\sim\mathcal{D}}\left[D_{\mathrm{KL}}(\bar{\pi}_{\omega}(\cdot|s,\omega)||\pi_{\theta}(\cdot|s,\omega))\right],$ (8)
where $\mathcal{D}$ is a replay buffer. By minimizing the above loss, we can combine policies for each preference into a single universal policy. After the training is completed, users can utilize various policies by providing preferences to the trained universal policy without additional training.
4.4 Convergence Analysis
In this section, we analyze the convergence properties of the proposed method. Specifically, it can be shown that CoMOGA achieves CP optimal convergence with minor modifications if $H$ is the Fisher information matrix. To this end, we first introduce the necessary conditions for the policy gradient method to ensure CP optimal convergence. We then show that the modified CoMOGA satisfies these necessary conditions. Before that, we first introduce the following assumption:
Assumption 4.1 (Slater’s condition).
There exists a feasible policy $\pi_{f}$ such that $J_{C_{k}}(\pi_{f})\leq d_{k}-\eta$ $\forall k$, where $\eta>0$ is a positive number.
The Slater’s condition is a common assumption in safe RL to ensure optimal convergence [20, 6, 11]. We also assume that the state and action spaces are finite. Now, we introduce a generalized version of the policy gradient method that guarantees optimal convergence.
@@ 4.4 Convergence Analysis
Assume that sequences $\nu^{a}_{t,i}$, $\nu^{b}_{t,i}$, $\lambda^{a}_{t,k}$, $\lambda^{b}_{t,k}\in[0,\lambda_{\mathrm{max}}]$ are given for all $i$ and $k$, where $\sum_{k}\lambda^{b}_{t,k}=1$, $\lambda^{b}_{t,k}(J_{C_{k}}(\theta_{t})-d_{k})\geq 0$, $\sum_{i}\nu^{a}_{t,i}=1$, and $\nu^{a}_{t,i}$ converges to a specific point $\bar{\nu}^{a}_{i}$. Then, if a policy is updated according to the following rule, it converges to a CP optimal policy of (1):
$\theta_{t+1}=\theta_{t}+\begin{cases}\alpha_{t}F^{\dagger}(\theta_{t})\left(\sum_{i=1}^{N}\nu^{a}_{t,i}g_{i}-\alpha_{t}\sum_{k=1}^{M}\lambda^{a}_{t,k}b_{k}\right)&\text{if}\;J_{C_{k}}(\theta_{t})\leq d_{k}\;\forall k,\\ \alpha_{t}F^{\dagger}(\theta_{t})\left(\alpha_{t}\sum_{i=1}^{N}\nu^{b}_{t,i}g_{i}-\sum_{k=1}^{M}\lambda^{b}_{t,k}b_{k}\right)&\text{otherwise},\\ \end{cases}$ (9)
where $g_{i}=\nabla J_{R_{i}}(\theta_{t})$, $b_{k}=\nabla J_{C_{k}}(\theta_{t})$, $\alpha_{t}$ is a step size satisfying Robbins-Monro condition [28], and $F^{\dagger}$ represents the pseudo-inverse of the Fisher information matrix.
The proof is provided in Appendix B.2. The sequences $\nu^{a}_{t,i}$, $\nu^{b}_{t,i}$, $\lambda^{a}_{t,k}$, and $\lambda^{b}_{t,k}$ in Theorem 4.2 represent the weights assigned to the gradients of the objective and constraint functions at each update step, and a new CMORL algorithm can be developed based on how the sequences are set. Similarly, we will show that CoMOGA can be formulated as (9) by identifying these sequences with minor modifications. Since the quadratic programming holds strong duality, we can express the solution of (5) as $\bar{g}_{\omega}^{\mathrm{ag}}=H^{-1}(\sum_{i=1}^{N}\nu^{*}_{i}g_{i}-\sum_{k=1}^{M}\lambda^{*}_{k}H^{-1}b_{k})$, where $\nu^{*}_{i}$ and $\lambda_{k}^{*}$ are the optimal Lagrange multipliers of the following dual problem:
$\nu^{*},\lambda^{*}=\underset{\nu\geq 0,\lambda\geq 0}{\mathrm{argmax}}$ $-\frac{1}{2}\left(\sum\nolimits_{i}\nu_{i}g_{i}-\sum\nolimits_{k}\lambda_{k}b_{k}\right)^{T}H^{-1}\left(\sum\nolimits_{i}\nu_{i}g_{i}-\sum\nolimits_{k}\lambda_{k}b_{k}\right)+\sum\nolimits_{i}\nu_{i}\omega_{i}e_{i}$
$+\sum\nolimits_{k}\lambda_{k}(J_{C_{k}}(\theta_{\mathrm{old}})-d_{k}).$
In case of the dual problem has no solution, we set $\nu_{i}^{*}$ and $\lambda_{k}^{*}$ to be $\bar{g}_{\omega}^{\mathrm{ag}}=0$. For the solution of (6), the same process is applied to achieve $\lambda_{k}^{*}$. Then, the CoMOGA is modified as follows:
$\bar{g}_{\omega}^{\mathrm{ag}}=H^{-1}\cdot\begin{cases}\sum_{i=1}^{N}\frac{\nu_{i}^{*}}{\sum_{j}\nu_{j}^{*}}g_{i}-\epsilon_{t}\sum_{k=1}^{M}\min(\frac{\lambda_{k}^{*}}{\epsilon_{t}\sum_{j}\nu_{j}^{*}},\lambda_{\mathrm{max}})b_{k}&\!\!\text{if}\;J_{C_{k}}(\theta_{t})\leq d_{k}\;\forall k,\\ -\sum_{k=1}^{M}(\lambda_{k}^{*}/(\sum_{j}\lambda_{j}^{*}))b_{k}&\!\!\text{otherwise},\\ \end{cases}$ (10)
$g_{\omega}^{\mathrm{ag}}=\epsilon_{t}\bar{g}_{\omega}^{\mathrm{ag}}/\min(\max(||\bar{g}_{\omega}^{\mathrm{ag}}||_{H},g_{\mathrm{min}}),g_{\mathrm{max}}),$
where $g_{\mathrm{min}}$, $g_{\mathrm{max}}$, $\lambda_{\mathrm{max}}\in\mathbb{R}_{\geq 0}$ are hyperparameters, and $\epsilon_{t}$ is the local region size satisfying Robbins-Monro condition. Finally, we can show optimal convergence of the modified CoMOGA.
@@
If $H$ is the Fisher information matrix and a policy $\pi_{\theta_{t}}$ is updated as $\theta_{t+1}=\theta_{t}+g_{\omega}^{\mathrm{ag}}$, where $g_{\omega}^{\mathrm{ag}}$ is defined in (10), it converges to a CP optimal policy of (1).
Theorem 4.3 can be proved by identifying sequences $\nu_{t,i}^{a}$, $\nu_{t,k}^{b}$, $\lambda_{t,i}^{a}$, $\lambda_{t,k}^{b}$ that satisfy the conditions mentioned in Theorem 4.2, and details are provided in Appendix B.3.
@@ 5 Practical Implementation
We need to calculate the gradient of the objective and constraint functions to perform CoMOGA and the universal policy update. To do that, we use reward and cost critics $V_{R,\psi_{R}}^{\pi}$, $V_{C,\psi_{C}}^{\pi}$, where $\psi_{R}$ and $\psi_{C}$ are the critic network parameters, to estimate the objectives and constraint function. The critics are trained by minimizing the following loss functions:
$\mathcal{L}(\psi_{R})$ $:=\underset{(s,a,s^{\prime},\omega)\sim\mathcal{D}}{\mathbb{E}}\left[\sum\nolimits_{i=1}^{N}(Y_{R_{i}}-V^{\pi}_{R_{i},\psi_{R}}(s,a,\omega))^{2}\right],$ (11)
where the target $Y_{R_{i}}$ can be calculated either as $R_{i}(s,a,s^{\prime})+\gamma V^{\pi}_{R_{i},\psi_{R}}(s^{\prime},a^{\prime},\omega)$ with $a^{\prime}\sim\pi(\cdot|s^{\prime},\omega)$ or using Retrace($\lambda$) [26], and the same process is applied to the cost critics. Given a preference $\omega$, the objective and constraint functions can be estimated as follows:
$J_{R_{i}}(\theta)$ $\approx\underset{\begin{subarray}{c}s\sim\mathcal{D},\\ a\sim\pi_{\theta}(\cdot|s,\omega)\end{subarray}}{\mathbb{E}}\left[V_{R_{i},\psi_{R}}^{\pi}(s,a,\omega)\right],\;J_{C_{k}}(\theta)\approx\underset{\begin{subarray}{c}s\sim\mathcal{D},\\ a\sim\pi_{\theta}(\cdot|s,\omega)\end{subarray}}{\mathbb{E}}\left[V_{C_{k},\psi_{C}}^{\pi}(s,a,\omega)\right].$ (12)
Then, we can compute the policy gradients using the critic networks through the reparameterization trick, as done in the SAC paper [15]. Finally, the proposed method is summarized in Algorithm 1.
Algorithm 1 Policy Update Using CoMOGA
@@
        Collect a trajectory $\tau=\{(s_{t},a_{t},r_{t},c_{t},s_{t+1})\}_{t=1}^{T}$ by using $\pi_{\theta}(\cdot|\cdot,\omega)$ and store $(\tau,\omega)$ in $\mathcal{D}$.
@@
     Update reward and cost critic networks using Equation (11).
@@
        Sample $(\tau,\omega)\sim D$, and estimate the objectives and constraints using Equation (12).
        Calculate the aggregated gradient $g^{\mathrm{ag}}_{\omega}$ using Equation (10), and set $\theta_{\omega}=\theta_{\mathrm{old}}+g^{\mathrm{ag}}_{\omega}$.
        Store an intermediate policy $\bar{\pi}_{\omega}=\pi_{\theta_{\omega}}$.